Multiple choice: 1, DeGiaimo Co. has an operating leverage of 5.next years sales are expected to increase by 10%, then the companys operating income will increase by 50%. a. True b. False 2. the unit selling price is $40, the volume of sales is $3,000,000, sales at the break- even point amount to $2.500,000, and the maximum possible sales are $3,300,000, the margin of safety will be 12,500 units. a. True b. False 3. Iit the unit seling price is $50, the volume of sales is $450,000, sales at the break even point amount to $375,000, and the maximum possible sales are $$50,000, the margin of safety will be 2.000 units a. True b. False 4. If the volume of sales is $4,000,000 and sales at the break-even point amount to $3200,000, the margin of safety will be 20%. a. True b. False . Iif the volume of sales is $6,000,000 and sales at the break-even point amount to $4,800,000, the margin of safety will be 25%. a. True b. False 8 Iif the volume of sales is $6,000,000 and sales at the break-even point amount to $5,000,000, the margin of safety wil b@ 20%. a. True b. False 6. Cost behavior refers to the manner in which: a. a cost changes as the related activity changes b. a cost is allocated to products с.acost is used in setting seling pnces. d. a cost is estimated 7. Costs that remain constant on a per-unit level as the level of activity changes are caled: a. fixed costs. b. mixed costs c. opportunity costs. d. variable costs. 8. Which of the following graphs ilustrates the behavior of a total foxed cost within the specified relevant range?
media%2F39b%2F39b38980-6d28-4c22-9082-14
media%2F7da%2F7daa6470-d1a6-43ee-9be8-7f
media%2F35c%2F35c1b12f-cccf-4633-a057-41
media%2F568%2F56827986-a199-4416-a26f-b2
media%2F6d3%2F6d38f94d-52fd-48ba-b06f-08


media%2Fe9e%2Fe9e3f9d3-2ae9-4594-878a-33
media%2F9bf%2F9bf8d56b-ad83-4573-9ca2-2d
Answer #1

Answer 1: True, OL of 5 means 10% change of sales will lead to 50% increase in Operating Income.

Answer 2: True, Margin of Safety = (Actual Sales - Break Even point)/ Sales Price per unit

We get, ($3000000-$2500000)/$40 = 12500 units of MoS.

Answer 3: False, Following the above formula, we will get a Margin of Safety of 1500 units and not 2000 units.

Answer 4: True, In percentage terms MoS shall be 20% i.e. Sales - BEP = $4m - $3.2m = $800000 i.e. 20% of Sales

Answer 5: False, the margin of safety will be 20% of sales and not 25% of sales

Answer 6: False, the margin of safety will be 16.67% and not 20% of sales. Answer can be found = (Sales - BEP/ Sales)*100 = (1000000/6000000)*100 = 16.67%

Answer 7: Option A, Cost behavior means to the manner cost changes as related acitivity of an organization changes. This is basically an analysis of how cost reacts to various levels of activities of an organization with increase and fall in their operations.